OpenStudy (anonymous):

If y = (x + 3)2, and the value of x=5, what is the value?

OpenStudy (anonymous):

the answer is 16.. Just substitute x with 5 and work it out you will get y=2(5+3) which is y=2(8) y=16

OpenStudy (unklerhaukus):

\[y(x) = (x + 3)^2\\ y(5)=(5+3)^2\\\qquad=(5+3)\times(5+3)\\\qquad=\]

OpenStudy (anonymous):

Thankies So Much Guys!! I Appreciate you guys showing your work!

OpenStudy (anonymous):

y =(5+3)*2 y= (8)*2 y= 16